OpenStudy (anonymous):

if f(x)=x^2 -1. what is the equation for f^-1(x)

OpenStudy (anonymous):

this function is not one to one so doesn't have an inverse

OpenStudy (anonymous):

....f^-1(x)= +/- (squareroot all)x+1 ...f^-1(x)= +/- (squareroot all )x-1 ....f^-1(x)= +/- (squareroot x only) x+1 ....f^-1(x)= +/- (squareroot x only) x-1

OpenStudy (anonymous):

if these are you choices you must pick one and then tell your on line system that they are giving wrong math